OpenStudy (anonymous):

Monica's heart beat 225 times in 3 minutes. Jordy's heart beat 360 times in 5 minutes. Whose heart beat at the faster rate and what was that rate? A. Monica had the faster heart rate of 72 beats per minute. B. Monica had the faster heart rate of 75 beats per minute. C. Jordy had the faster heart rate of 72 beats per minute. D. Jordy had the faster heart rate of 75 beats per minute.

OpenStudy (luckyness):

\[\frac{ 225 }{ 3 }\]

OpenStudy (luckyness):

Solve for that ^

OpenStudy (zaynab123):

Monica because if you divide 225 by 3 you get 75 and divide 360 by 5 is 72

OpenStudy (luckyness):

It is B. Take monica: 225/3=75 And now take Jordy: 360/5=72 Monicas is higher.

OpenStudy (anonymous):

thanx